# It’s Not You, It’s Clipping: A Soft Trust-Region via Probability Smoothing for LLM RL

### v1

@@ Abstract
Training large language models (LLMs) with reinforcement learning (RL) methods such as PPO and GRPO commonly relies on ratio clipping to stabilise updates. While effective at preventing instability, clipping discards information and introduces gradient discontinuities. We propose Probability Smoothing Policy Optimisation (PSPO), which smooths the current policy’s probabilities toward the old (behaviour) policy before computing the importance ratio, analogous to label smoothing. Unlike clipping, PSPO preserves gradient signal, while interpolation toward the old policy creates a soft trust region that discourages large, destabilising updates, with formal guarantees.
We instantiate PSPO within GRPO (GR-PSPO) and fine-tune Qwen2.5-0.5B/1.5B on GSM8K, evaluating on GSM8K test and the cross-dataset generalisation on SVAMP, ASDiv, and MATH-500. Relative to unclipped GRPO (single iteration; no data reuse, ratio always = 1), GR-PSPO attains similar accuracy but produces clearer, more concise, and more logically coherent responses (LLM-as-Judge). Compared to clipped GRPO, GR-PSPO substantially improves performance in both the 0.5B and 1.5B models, with a boost of over 20% on GSM8K (39.7% vs. 17.6% for 0.5B, 59.4% vs. 37.8% for 1.5B).
@@ 1 Introduction
Reinforcement learning (RL) is now a central component of large language model (LLM) fine-tuning pipelines after supervised fine-tuning (SFT) (Ouyang et al., ). Proximal Policy Optimization (PPO; Schulman et al., 2017b) underpins systems such as WebGPT (Nakano et al., ), LLaMA-2 Chat (Touvron et al., 2023), and Sparrow (Glaese et al., ). Group Relative Policy Optimization (GRPO) adapts PPO for LLMs (Shao et al., 2024) and has been applied to mathematical reasoning tasks (Shao et al., 2024), alongside other RL approaches (Luong et al., 2024; Mitra & Ulukus, 2025; Luo et al., 2025; Zheng et al., 2025). A key challenge within policy optimisation is achieving a balance between learning speed and stability. Optimal theoretical options (Schulman et al., 2017a) lead to small step sizes, making convergence at best inefficient but often infeasible. Trust Region Policy Optimisation (TRPO) (Schulman et al., 2017a) constrains updates using the KL divergence, which allows for larger steps but is computationally inefficient. PPO provides an empirically stronger regularisation by using clipped probability ratios as a first-order approximation of the KL divergence, of which GRPO inherits.
However, ratio clipping has drawbacks, namely vanishing gradients when the policy ratio leaves the clip range. Additionally, clipping can miss better policies outside of the clipped policy space (Chen et al., 2022), especially in problems where greater exploration might be beneficial. Alternatives (KL early stopping (Sun et al., 2022), smooth transforms (Chen et al., 2022)) can be brittle or saturating, particularly in more complex settings. Some implementations of GRPO (Hugging Face, ) effectively avoid clipping by using a single pass over data; the importance sampling ratio is always 1, which essentially reverts the approach back to a vanilla policy gradient method, and as such, will typically require small steps and be sample inefficient.
@@
We propose Probability Smoothing Policy Optimisation (PSPO), as an alternative to clipping. Instead of truncating ratios, we smooth the current policy’s probabilities toward the old behaviour policy before computing the importance ratio. This is inspired by label smoothing in supervised learning. This smoothing reduces overconfidence in any single action while retaining informative gradients everywhere. Crucially, by interpolating with $\pi_{\theta_{\text{old}}}$, it acts as a soft trust region. We instantiate PSPO in GRPO (GR-PSPO) and evaluate on GSM8K (Cobbe et al., 2021), ASDiv (Miao et al., 2020), SVAMP (Patel et al., 2021), and MATH-500 (Lightman et al., 2023), training on Qwen2.5-0.5B/1.5B.
@@ 2 Probability Smoothing Policy Optimisation
Policy gradient methods optimise the expected reward by updating the policy $\pi_{\theta}$ with respect to sampled trajectories. To effectively reuse trajectories from an old policy $\pi_{\theta_{\text{old}}}$, any update is regularised using importance sampling. Importance sampling estimates how likely the (state $s_{t}$, action $a_{t}$) pair would occur given the current policy. In PPO (Schulman et al., 2017b) and GRPO Shao et al. (2024), this is approximated with a ratio of the current policy $\pi_{\theta}$ and the old, behaviour policy $\pi_{\theta_{\text{old}}}$ which generated the trajectory. This ratio is defined in equation  1,
@@
GRPO (Shao et al., 2024) is an adaptation of PPO, which removes the need for a critic model, reducing the amount of training resources and developed specifically for LLMs. GRPO samples a group $G$ of outputs $a$ for a given prompt $s\in S$, and uses the group scores $r$ as a baseline estimate to then calculate the advantage $\hat{A}$ using the relative rewards based on the current group baseline; $\hat{A}_{t,i}=R_{t,i}-\bar{R}_{t,i}$. GRPO includes the same clipping principle as PPO in its surrogate objective, although some default implementations suggest that using GRPO with only 1 iteration over the data gives comparable performance and negates the effect of clipping. GRPO aims to maximise the objective function:
$J^{\text{GRPO}}(\theta)=\mathbb{E}_{t}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\Bigg\{\min\Big(r_{t,i}(\theta)\hat{A}_{t,i},\,\text{clip}(r_{t,i}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}_{t,i}\Big)-\beta\mathbb{D}_{KL}[\pi_{\theta}\mid\pi_{ref}]\Bigg\}\Bigg],$ (2)
@@
To reduce overconfidence in any single action in a given state, we took inspiration from the label smoothing regularisation method used in supervised learning (Szegedy et al., 2016). Label smoothing has been shown to reduce overconfidence and improve the robustness of a model (Müller et al., ) (Goibert & Dohmatob, 2019). Label smoothing, equation  3, moves from one-hot encoded target distribution $\varphi(k\mid x)$ to soft targets $\tilde{\varphi}(k\mid x)$ that are a weighted average of the hard target distribution and another distribution, traditionally the uniform distribution $u(k)$ (Szegedy et al., 2016),  [^1] In the original label smoothing paper, $\varepsilon$ is used as the smoothing parameter, we use $\alpha$ to avoid confusion with the clipping range, often denoted as $\varepsilon$.
@@
where $\alpha\in[0,1]$ controls the smoothing strength. We apply (3) to the current policy probability in equation  4,
@@
For policy optimisation, updates should be within a trust region to enable larger step updates. Therefore, we decided to smooth towards the old behaviour policy, $q=\pi_{\theta_{\text{old}}}$, rather than the uniform distribution, so the smoothing behaves as a behaviour-anchored trust region.  Szegedy et al. (2016) noted that the deviation in the loss when using label smoothing compared to the loss otherwise could equivalently be captured by the KL divergence. This bolsters our intuition to smooth towards the old policy, and in doing so, we introduce an equivalent estimate of the KL-divergence with the smoothed action probability and create a soft trust region. The smoothed probability becomes equation  5,
@@
which given (5), becomes equation  7,
$\tilde{r}_{t}(\theta)=\frac{(1-\alpha)\pi_{\theta}(a_{t}\mid s_{t})+\alpha\cdot\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}=(1-\alpha)r_{t}+\alpha.$ (7)
@@
Lemma 1 (Total variation contraction).
@@
Corollary 1 (KL upper bounds shrink under smoothing).
@@
$\lambda P_{1}+(1-\lambda)P_{2}=(1-\alpha)\pi_{\theta}+\alpha\pi_{\theta_{\text{old}}}=\tilde{\pi_{\theta}},\quad\lambda Q_{1}+(1-\lambda)Q_{2}=(1-\alpha)\pi_{\theta_{\text{old}}}+\alpha\pi_{\theta_{\text{old}}}=\pi_{\theta_{\text{old}}}$
@@
Hence $\alpha$ directly sets a soft trust-region radius in both TV and (upper-bounded) KL.
Proposition 1 (Ratio contraction and non-vanishing slopes).
@@
Proposition 2 (Overconfidence regularisation).
@@
Proposition 3 (PSPO surrogate as a scaled policy gradient with implicit stability).
@@
$\mathcal{J}_{\text{PSPO}}(\theta)=\mathbb{E}_{a\sim\pi_{\theta_{\text{old}}}}\![\tilde{r}(a)\,A(a)]=(1-\alpha)\,\mathbb{E}_{a\sim\pi_{\theta}}\![A(a)]+\alpha\,\mathbb{E}_{a\sim\pi_{\theta_{\text{old}}}}\![A(a)].$
using equation 1 and the change of measure formula. Only the first term depends on $\theta$, and using policy gradient theorem, $\nabla_{\theta}\mathcal{J}_{\text{PSPO}{}}=(1-\alpha)\,\mathbb{E}_{a\sim\pi_{\theta}}\!\big[\nabla_{\theta}\log\pi_{\theta}(a\mid s)\,A(a)\big].$ Hence, PSPO is the on-policy gradient scaled by $(1-\alpha)$, while the policy itself is mixed with $\pi_{\theta_{\text{old}}}$ (Lemma 1), implicitly controlling divergence without an explicit KL term (which we set $\beta{=}0$ in our GRPO runs).
@@
We demonstrate how PSPO can apply to GRPO, to produce GR-PSPO which changes (2) to equation  8,
$J^{\text{\text{GR-\text{PSPO}}{}}}(\theta)=\mathbb{E}_{t}\Bigg[\frac{1}{G}\sum_{i=1}^{G}(\tilde{r}_{t,i}(\theta)\hat{A}_{t,i})-\beta\cdot\mathbb{D}_{KL}[\pi_{\theta}\mid\pi_{ref}]\Bigg].$ (8)

### v2

@@ Abstract
Training large language models (LLMs) with reinforcement learning (RL) methods such as PPO and GRPO commonly relies on ratio clipping to stabilise updates. While effective at preventing instability, clipping discards information, introduces gradient discontinuities and can prevent exploration of better policies. Inspired by label smoothing, we propose Probability Smoothing Policy Optimisation (PSPO). PSPO smooths current policy probabilities toward the behaviour policy before computing importance ratios, creating a soft trust region that preserves gradients while preventing destabilising updates. Unlike prior soft clipping approaches that use sigmoid-based transformations which can suffer from vanishing gradients and saturation, our method uses a linear interpolation, providing simpler and more robust gradient preservation.
Empirically, GR-PSPO outperforms clipping and sigmoid-based alternatives on mathematical reasoning benchmarks when refining models with prior domain knowledge, achieving an accuracy of 79.9% on GSM8K and 59.6% on MATH for Qwen2-Math-1.5B.
@@ 1 Introduction
Reinforcement learning (RL) is now a central component of large language model (LLM) fine-tuning. RL is commonly used after supervised fine-tuning (SFT) (Ouyang et al., 2022), where the aim is to refine an existing policy whilst preserving prior knowledge. In this setting, the main challenge is to make improvements without degrading existing knowledge, as unstable, destructive updates can waste expensive training. RL post-training has shown enhanced reasoning and can maintain context in long and complex conversations. This is especially important with the prevalence of Small Language Models, that due to their size are energy efficient and easier to control in a workflow as they can be trained and maintained by smaller companies but where these models in particular struggle to provide outputs that are understandable by a human user.
To stabilise updates, policy optimisation methods rely on constraints that limit deviation from a reference policy. In practice, this requires a balance between learning speed and stability, as overly conservative updates can make learning inefficient or infeasible. Trust Region Policy Optimisation (TRPO) (Schulman et al., 2015) constrains updates using the KL divergence, which allows for larger steps but is still computationally inefficient. Proximal Policy Optimization (PPO) (Schulman et al., 2017) provides an empirically stronger regularisation by using clipped probability ratios as a first-order approximation of the KL divergence. PPO clips the importance sampling ratio to effectively reduce the step size when the policy is outside the trust region, and underpins systems such as WebGPT (Nakano et al., 2021), LLaMA-2 Chat (Touvron et al., 2023), and Sparrow (Glaese et al., 2022).
This is extended in Group Relative Policy Optimization (GRPO) which is an adaptation of PPO specifically for LLMs  (Shao et al., 2024), inheriting the clipped probability, and has been applied to mathematical reasoning tasks (Shao et al., 2024), alongside other RL approaches (Luong et al., 2024; Mitra and Ulukus, 2025; Luo et al., 2025; Zheng et al., 2025). Some implementations of GRPO (Hugging Face, ) default to using a single pass over data, effectively avoiding clipping by only using on-policy data. This leads to the importance sampling ratio always being 1, which essentially reverts the approach back to a vanilla policy gradient method, and as such, will typically require small steps and be sample inefficient.
However, ratio clipping has drawbacks, namely vanishing gradients when the policy ratio leaves the clip range. Additionally, clipping can miss better policies outside of the clipped policy space (Chen et al., 2023), especially in problems where greater exploration might be beneficial. Alternatives to clipping have therefore emerged, predominantly focusing on smooth transforms to remove the brittleness of hard clipping. P3O (Chen et al., 2023) introduced sigmoid-based soft clipping through its SCOPIC objective to enable exploration beyond the trust region defined by typical hard clip boundaries. P3O was empirically shown to perform better than PPO on five deep RL gym environments, but has not been applied to GRPO or used for LLM fine-tuning prior to this work. Sigmoid methods can, however, lead to saturation of values at extreme ratios, reintroducing vanishing gradients in precisely the regions where clipping also fails. Recent work has addressed the asymmetric behaviour of positive versus negative advantages in LLM fine-tuning specifically, Tapered Off-Policy REINFORCE (TOPR) (Roux and Bellemare, 2025) employs asymmetric importance sampling to differentially weight positive and negative advantages showing better performance on GSM8K compared with PPO, DPO and REINFORCE on 8B parameter models. Soft Adaptive Policy Optimization (SAPO) (Gao et al., 2025) combines sigmoid soft gating with asymmetric control, similar to ideas from both P3O and TOPR, showing an increase in performance compared with GRPO and GSPO on mathematical tasks on 30B parameter models. Asymmetric methods require at least one additional parameter to be tuned, and TOPR’s performance seems to be quite dependent on the ratio of positive to negative samples in the training, which is not always known a priori or easy to control especially given that it will change across training. Other approaches include KL early stopping (Sun et al., 2022), though these can be brittle or saturating, particularly in complex settings.
@@
We propose Probability Smoothing Policy Optimisation (PSPO), as an alternative to clipping. Instead of truncating ratios, we smooth the current policy’s probabilities toward the old behaviour policy before computing the importance ratio. This is inspired by label smoothing in supervised learning. This smoothing reduces overconfidence in any single action while retaining informative gradients everywhere.
We empirically evaluate PSPO by instantiating it within GRPO (Shao et al., 2024) (GR-PSPO) across three model types: Qwen2.5-0.5B, Qwen2.5-1.5B, and Qwen2-Math-1.5B. This experimental design allows us to test our hypothesis that PSPO’s advantages emerge when stability is required, such as with models with prior domain knowledge (the Math model).
@@ 2 Probability Smoothing Policy Optimisation
Policy gradient methods optimise the expected reward by updating the policy $\pi_{\theta}$ with respect to sampled trajectories. To effectively reuse trajectories from an old policy $\pi_{\theta_{\text{old}}}$, any update is regularised using importance sampling. Importance sampling estimates how likely the (state $s_{t}$, action $a_{t}$) pair would occur given the current policy. In PPO (Schulman et al., 2017) and GRPO Shao et al. (2024), this is approximated with a ratio of the current policy $\pi_{\theta}$ and the old, behaviour policy $\pi_{\theta_{\text{old}}}$ which generated the trajectory. This ratio is defined in equation  1,
@@
GRPO is an adaptation of PPO,  (Shao et al., 2024), which removes the need for a critic model, reducing the amount of training resources and developed specifically for LLMs. GRPO samples a group $G$ of outputs $a$ for a given prompt $s\in S$, and uses the group scores $r$ as a baseline estimate to then calculate the advantage $\hat{A}$ using the relative rewards based on the current group baseline; $\hat{A}_{t,i}=R_{t,i}-\bar{R}_{t,i}$. GRPO includes the same clipping principle as PPO in its surrogate objective, although some default implementations suggest that using GRPO with only 1 iteration over the data gives comparable performance and negates the effect of clipping. GRPO aims to maximise the objective function:
$\begin{split}J^{\text{GRPO}}(\theta)=\mathbb{E}_{t}\Bigg[\frac{1}{G}\sum_{i=1}^{G}\Bigg\{\min\Big(r_{t,i}(\theta)\hat{A}_{t,i},\,\text{clip}(r_{t,i}(\theta),1-\varepsilon,1+\varepsilon)\hat{A}_{t,i}\Big)-\beta\mathbb{D}_{KL}[\pi_{\theta}\mid\pi_{ref}]\Bigg\}\Bigg],\end{split}$ (2)
@@
To reduce overconfidence in any single action in a given state, we took inspiration from the label smoothing regularisation method used in supervised learning (Szegedy et al., 2016). Label smoothing has been shown to reduce overconfidence and improve the robustness of a model (Müller et al., 2019) (Goibert and Dohmatob, 2019). Label smoothing, equation  3, moves from one-hot encoded target distribution $\varphi(k\mid x)$ to soft targets $\tilde{\varphi}(k\mid x)$ that are a weighted average of the hard target distribution and another distribution, traditionally the uniform distribution $u(k)$ (Szegedy et al., 2016),  [^1] In the original label smoothing paper, $\varepsilon$ is used as the smoothing parameter, we use $\alpha$ to avoid confusion with the clipping range, often denoted as $\varepsilon$.
@@
where $\alpha\in[0,1]$ controls the smoothing strength.
We apply (3) to the current policy probability in equation  4,
@@
For policy optimisation, updates should be within a trust region to enable larger step updates. Therefore, we decided to smooth towards the old behaviour policy, $q=\pi_{\theta_{\text{old}}}$, rather than the uniform distribution, so the smoothing behaves as a behaviour-anchored trust region.  Szegedy et al. (2016) noted that the change in the loss when using label smoothing could be equivalently considered as an estimate of the KL divergence. Given that PPO used the probability ratio clipping as a first-order estimate of KL-divergence for regularisation, this bolsters our intuition to smooth towards the old policy, and by doing so, we introduce an equivalent estimate of the KL-divergence with the smoothed action probability and create a soft trust region. The smoothed probability becomes equation  5,
@@
which given (5), becomes equation 7,
$\begin{split}\tilde{r}_{t}(\theta)=\frac{(1-\alpha)\pi_{\theta}(a_{t}\mid s_{t})+\alpha\cdot\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}\mid s_{t})}\\ =(1-\alpha)r_{t}+\alpha.\end{split}$ (7)
Having derived the smoothed ratio (Eq. 7) we now establish formally that this simple linear transformation provides the properties we intuitively assumed: (i) contraction toward the behaviour policy in both total variation and KL divergence, ensuring updates remain within a trust region; (ii) preservation of non-zero gradients everywhere, avoiding the learning signal loss caused by clipping; and (iii) regularisation against overconfidence in any single action. We state these results as formal propositions.
@@
Lemma 2.1 (Total variation (TV) contraction).
@@
This means $\alpha$ directly controls the maximum distance the smoothed policy can drift from the behaviour policy: setting $\alpha=0.1$ guarantees the smoothed policy is at most 90% as far from $\pi_{\theta_{\text{old}}}$ as the unsmoothed policy would be, in total variation distance. This provides a tunable constraint on policy updates without hard truncation.
Corollary 2.2 (KL upper bounds shrink under smoothing).
@@
$\begin{split}\lambda P_{1}+(1-\lambda)P_{2}=(1-\alpha)\pi_{\theta}+\alpha\pi_{\theta_{\text{old}}}=\tilde{\pi_{\theta}},\\ \lambda Q_{1}+(1-\lambda)Q_{2}=(1-\alpha)\pi_{\theta_{\text{old}}}+\alpha\pi_{\theta_{\text{old}}}=\pi_{\theta_{\text{old}}}\end{split}$
@@
Hence $\alpha$ directly sets a soft trust-region radius in both total variation (TV) and (upper-bounded) KL.
This establishes that probability smoothing induces an implicit KL penalty, which can be viewed as a soft trust-region with radius of $\alpha$.
Proposition 2.3 (Ratio contraction and non-vanishing slopes).
@@
This is the key advantage over clipping. With PPO/GRPO clipping, when $r>1+\varepsilon$ (for positive advantages) or $r<1-\varepsilon$ (for negative advantages), the gradient is exactly zero, the model receives no learning signal for these samples. PSPO instead scales the gradient by $(1-\alpha)$, which is non-zero everywhere. This means PSPO continues to learn from samples that clipping would ignore, while the contraction still prevents destabilising large updates.
Proposition 2.4 (Overconfidence regularisation).
@@
This property prevents the policy from over fitting and becoming overly confident in any single action. When the current policy assigns higher probability to an action than the behaviour policy did, smoothing pulls this probability down. This is particularly relevant for language generation, where multiple tokens may be equally valid continuations.
Proposition 2.5 (PSPO surrogate as a scaled policy gradient with implicit stability).
@@
$\begin{split}\mathcal{J}_{\text{PSPO}}(\theta)=\mathbb{E}_{a\sim\pi_{\theta_{\text{old}}}}\![\tilde{r}(a)\,A(a)]\\ =(1-\alpha)\,\mathbb{E}_{a\sim\pi_{\theta}}\![A(a)]+\alpha\,\mathbb{E}_{a\sim\pi_{\theta_{\text{old}}}}\![A(a)].\end{split}$
using equation 1 and the change of measure formula. Only the first term depends on $\theta$, and using policy gradient theorem, $\nabla_{\theta}\mathcal{J}_{\text{PSPO}{}}=(1-\alpha)\,\mathbb{E}_{a\sim\pi_{\theta}}\!\big[\nabla_{\theta}\log\pi_{\theta}(a\mid s)\,A(a)\big].$ Hence, PSPO is the on-policy gradient scaled by $(1-\alpha)$, while the policy itself is mixed with $\pi_{\theta_{\text{old}}}$ (Lemma 2.1), implicitly controlling divergence without an explicit KL term (which we set $\beta{=}0$ in our GRPO runs).
This shows that PSPO can be seen as performing standard on-policy gradient descent, but with two modifications: the gradient is scaled by $(1-\alpha)$, slowing learning, while the policy used to compute advantages is mixed with $\pi_{\theta_{\text{old}}}$. Together, these effects stabilise learning by anchoring updates to recent behaviour.
Proposition 2.5 shows that PSPO scales the policy gradient by a factor of $1-\alpha$. While this may resemble a simple reduction in the learning rate, PSPO is not equivalent to learning-rate scaling alone. The scaling arises from probability mixing with the behaviour policy, which simultaneously contracts the policy distribution toward $\pi_{\theta_{\text{old}}}{}$. As a result, PSPO adaptively alters the update magnitudes based on the model’s own prior probabilities, rather than applying a uniform step-size reduction. In contrast, clipping also implicitly rescales gradients, but does so discontinuously by truncating updates past a fixed ratio threshold.
@@
Selecting $\alpha$.
The smoothing parameter $\alpha\in[0,1]$ controls the strength of the trust region by determining the fraction of the old policy to mix with the current policy. The larger the value of $\alpha$, the stronger the regularisation is, and the ratio is contracted more toward 1, and the KL upper bounds shrink (Corollary 2.2). A larger $\alpha$ is analogous to a smaller clipping range $\varepsilon$ in traditional GRPO. In practice, we have found that $\alpha\in[0.1,0.4]$ works well across our experiments, comparable to typical clipping ranges used in PPO/GRPO. Importantly, $\alpha$ replaces $\varepsilon$ as the trust region hyperparameter: both methods require tuning exactly one scalar controlling the constraint strength, so PSPO introduces no additional hyperparameter burden.
@@
We demonstrate how PSPO can apply to GRPO, denoted as GR-PSPO which changes (2) to equation  8,
$J^{\text{GR-PSPO}}(\theta)=\mathbb{E}_{t}\Bigg[\frac{1}{G}\sum_{i=1}^{G}(\tilde{r}_{t,i}(\theta)\hat{A}_{t,i})-\beta\cdot\mathbb{D}_{KL}[\pi_{\theta}\mid\pi_{ref}]\Bigg].$ (8)